OpenStudy (kenshin):

Please help me understand the following step in arriving at the answer to this question: \[x \sin (\pi*x) = \int\limits_{0}^{x^2} f(t) dt, find f(4);\] The answer has this step: differentiate both sides and get: \sin (\pi*x) + \pi*x*\cos(\pi*x) = 2*x*f(x^2)\] How did the upper limit become 2x here from chainrule?

OpenStudy (kenshin):

\[\sin (\pi*x) + \pi*x*\cos(\pi*x) = 2xf(x^2) this is the step i'm \not getting, where did the 2xf(x^2) come from, the answer said \it was from chainrule and FTC1\]